Just need simple explanation for understanding thank you
If residents of a country, collectively, want to save more than businesses want to invest, how would this affect its macro-economy? Explain. 6. llects in tax (or other) If the government of a country spends more than it co revenue, it is running a 7. What is inflation? Deflation? How do inflation and deflation affect the distribution of income and wealth in a country? 8. What is the difference between a capitalist and socialist (also known as communi economic system? 9.

5. The consumption function for a country is C-0.85Y. Explain what this means (in words). What does 0.85 represent? What is the savings rate in this country?
Answer #1

Answer:

Question 5)

giving C=0.85Y

Here 0.85 is represent Marginal Propensity to consume(MPC).

The marginal propensity to consume (MPC) is the proportion of an aggregate raise in salary that a consumer spends on the consumption of goods and services.

Saving rate=1-MPC=1-0.85=0.15